OpenStudy (anonymous):

Suppose your bedroom is 18 ft long, 15 ft wide, and the distance from floor to ceiling is 8 ft, 6 in. You need to know the volume of the room in metric units for some scientific calculations. (a) What is the room’s volume in cubic meters? In liters? (b) What is the mass of air in the room in kilograms? In pounds? (Assume the density of air is 1.2 g/L and that the room is empty of furniture.)

OpenStudy (anonymous):

do you know the volume in cubic feet?

OpenStudy (anonymous):

or cubic inches

OpenStudy (anonymous):

I think its 15.003 cubic feet

OpenStudy (anonymous):

2295 ft^3 \[2295 ft^3 \times \frac{ 0.3048^3 m^3 }{ 1 ft^3 } = 699.5\]

OpenStudy (anonymous):

m^3

OpenStudy (anonymous):

1m^3 = 1000L density = mass/volume so mass = density*volume